# Visual Hallucinations of Multi-modal Large Language Models

### v1

@@ Appendix A Generated Text Descriptions of VH Modes
Appendix B Prompts Used in VHTest
@@
Prompt for text description generation using a successful initial VH instance:
[image] Question: [question] Multi-modal LLM (MLLM)’s response: [testing MLLM’s hallucinated response] Detected hallucination in the response: [detected hallucination in the response] Focus on the following elements: image, question, MLLM’s response, and the detected hallucination in the response. I’m trying to identify visual hallucinations in the MLLM associated with its visual process. Through the specific example, are there any general types of hallucination modes you notice the MLLM makes, or any visual features that MLLM fails to encode, ultimately leading to the errors in MLLM’s response? Try to give hallucination modes that are specific enough that someone could enable consistent reproduction of images and corresponding questions. Please try to include as many general hallucination modes as possible. These hallucination modes will be used later to generate images or videos. In your hallucination modes, please clearly explain why the this hallucination mode would cause difficulties for the vision encoder of MLLM to understand images related to this hallucination mode. I will further use this reason to precisely generate images that match the hallucination mode and are able to mislead the MLLM. Please encapsulate the essence of the examples provided, summarize as many as possible and stick to the examples.
@@
Prompt for text description generation using an unsuccessful initial VH instance:
[image] Question: [question] A hypothetical response from the multi-modal LLM (MLLM): [a hypothetical hallucinated response] Detected hallucination in the hypothetical response: [detected hallucination in the hypothetical response] Focus on the following elements: image, question, MLLM’s hypothetical response, and detected hallucination in the hypothetical response. I’m trying to identify visual hallucinations in the MLLM associated with its visual process. Through the specific example, are there any general types of hallucination modes you notice the MLLM makes, or any visual features that MLLM fails to encode, ultimately leading to the errors in MLLM’s response? Try to give hallucination modes that are specific enough that someone could enable consistent reproduction of images and corresponding questions. Please try to include as many general hallucination modes as possible. These hallucination modes will be used later to generate images or videos. In your hallucination modes, please clearly explain why the this hallucination mode would cause difficulties for the vision encoder of MLLM to understand images related to this hallucination mode. I will further use this reason to precisely generate images that match the hallucination mode and are able to mislead the MLLM. Please encapsulate the essence of the examples provided, summarize as many as possible and stick to the examples.
@@
Prompt for text description integration:
I want you to focus on the [VH mode]: [the definition of VH mode]. Try to summarize the [VH mode] text description in ONE paragraph less than 200 words to explain the causes of the hallucination based on the information below. [N text descriptions]
@@
Prompt for VH image generation:
Generate an image that reflects the given hallucination mode. After your image generation, we will manually generate questions based on the image. Your MAIN GOAL is to ensure that the image aligns with the hallucination mode well so that querying the multi-modal LLM with the combination of manual questions and the image can effectively induce hallucinations. You will be evaluated on how well you actually perform. The generated image should ideally align with the hallucination mode, but there’s room for creativity. Be both creative and cautious. You can try to create image with different scenes and objects that align with the hallucination mode. Moreover, when you generate images, remember you are a very clever expert in exploiting the hallucination mode. For future debugging purposes, ensure that the generated images with our manual questions MUST cause multi-modal LLMs to provide incorrect responses. Hallucination Mode: [Text description of a VH mode]
@@
Prompt for an LLM to generate prompts for text-to-image generative models:
Generate an image that reflects the given hallucination mode. After your image generation, we will manually generate questions based on the image. Your MAIN GOAL is to ensure that the image aligns with the hallucination mode well so that querying the multi-modal LLM with the combination of manual questions and the image can effectively induce hallucinations. You will be evaluated on how well you actually perform. The generated image should ideally align with the hallucination mode, but there’s room for creativity. Be both creative and cautious. You can try to create image with different scenes and objects that align with the hallucination mode. Moreover, when you generate images, remember you are a very clever expert in exploiting the hallucination mode. For future debugging purposes, ensure that the generated images with our manual questions MUST cause multi-modal LLMs to provide incorrect responses. Hallucination Mode: [Text description of a VH mode] As an LLM, DO NOT generate images, but generate prompts for text-to-image generative models aligning with the hallucination mode. I will use your generated prompts for text-to-image generative models later.
@@
Appendix C Question Templates
@@
1. Could you describe [object] in the picture? 2. Describe the [property] of [object] in the picture.
Question templates for shape VH mode
@@
1. Describe the shape(s) of [object] in the picture. 2. In the photo, is [object] depicted with a conventional, typical shape?
Question templates for color VH mode
@@
1. Describe the color(s) of [object] in the picture. 2. Describe [n] main color(s) of [object] in the picture. 3. Describe all the colors of [object] in the picture. Do not miss any color. 4. Describe the original color of [object] in the picture.
Question templates for orientation VH mode
@@
1. From the camera/viewer’s perspective, please choose the direction that is most proximate to the orientation in which [object] is facing in the picture. [Left, 45 degrees leftwards away from the camera/viewer, Away from the camera/viewer, 45 degrees rightwards away from the camera/viewer, Right, 45 degrees rightwards towards the camera/viewer, Towards the camera/viewer, 45 degrees leftwards towards the camera/viewer] 2. From the camera/viewer’s perspective, please choose the direction that is most proximate to the orientation in which [object] is facing in the picture. [Left, Left and upward within the pictorial plane, Upward within the pictorial plane, Right and upward within the pictorial plane, Right, Right and downward within the pictorial plane, Downward within the pictorial plane, Left and downward within the pictorial plane] 3. From the camera/viewer’s perspective, please choose the direction that is most proximate to the orientation in which [object] is facing in the picture. [Away from the camera/viewer, Towards the camera/viewer]
Question templates for OCR VH mode
@@
1. Read the (English) character(s)/word(s)/ caption [at WHERE] in the picture.
Question templates for size VH mode
@@
1. Could you identify and locate [object] with the largest/smallest size in the picture? 2. Could you identify and locate [object] with the tallest/shortest height in the picture? 3. Could you identify and locate [object] with the second largest/smallest size in the picture? 4. Could you identify and locate [object] with the largest/smallest and the second largest/smallest size in the picture? 5. Could you identify and locate [object] with the tallest/shortest and the second tallest/shortest height in the picture? 6. Could you identify and locate the 2/3/4 largest/smallest [object] in the picture? 7. Which one is larger/smaller/taller/shorter, [object A] or [object B]? 8. What is the largest [object] in the picture?
Question templates for position VH mode
@@
1. Is [object A] to the left/right of/on the top of/being placed on/above/under/inside/outside [object B]? 2. Which one is on the left/right/top, [object A] or [object B]? 3. Which one is closer to/further from the camera or viewer perspective, [object A] or [object B]? 4. Which one is positioned under/positioned above/closer to [reference object], [object A] or [object B]? 5. Regardless of the positional relationship of the actual scene in reality, from the camera’s perspective in the photo, is [object A] positioned above/below [object B]? 6. Is [object A] on and touching [object B]?
Question templates for counting VH mode
Appendix D Special Cases in Step III
@@
Prompt to find names of non-existent objects in a VH image in the existence VH mode
[image] First, list all the names of objects in the picture. And then return some names of objects according to Requirement 1 or Requirement 2: Requirement 1: Objects you associate with the scene that should be there but are not actually there in the picture. Requirement 2: Objects that look similar to an object in the picture but do not actually exist in the picture.
@@
Appendix E Details of Fine-tuning Experiments
E.1 Hyperparamters
@@
E.2 Pre-processing of the Training Set
As mentioned in [14], the short-form data overfit an MLLM behaviorally to short-form answers. We follow this work to do pre-processing on both counting and position VH modes. To be more specific, for counting VH mode instances, we use the prompt below in Figure 14 for an LLM (e.g., GPT-4) to modify every number reference answer into a reference sentence. As for position VH mode instances, we add the sentence “Answer the question using a single word or phrase.” after all the questions in every position VH instance.
Prompt for reference answer modification in counting VH mode

### v2

@@ Appendix A Generated Text Descriptions of VH Modes
Appendix B Additional Results
B.1 Inter-annotator Agreement
Because annotating the whole benchmarks relies on heavy human labor, we randomly select 100 VH instances from our YNQ benchmark to measure the inter-annotator agreement. Specifically, four independent human annotators label these 100 VH instances. Based on these labeling results, we report Fleiss’ Kappa ($\kappa$) to indicate the inter-annotator agreement. We obtain $\kappa=0.958$ ($0.81\leq\kappa\leq$1.00), showing a very high degree of agreement.
B.2 More MLLMs Evaluation on VHTest
We evaluate 4 additional MLLMs on the YNQ benchmark: the Gemini-pro-vision Gemini (March 2024), ShareGPT4V-13b Chen et al. (2023b), InstructBLIP-13b Dai et al. (2023), and Qwen-VL-Chat-7b Bai et al. (2023), as shown in Table 7. Together with Table 3, GPT-4V shows the highest average accuracy of 0.611 among these 7 MLLMs.
B.3 Qualitative Comparison Results
Figure 8: Qualitative comparison results on mitigating visual hallucinations of LLaVA-1.5 via fine-tuning.
In Figure 8, we show some visual question answering examples using LLaVA-1.5 before and after fine-tuning on our benchmark. For example, before fine-tuning, LLaVA-1.5 incorrectly identifies the shape of a banana, as shown at the top row; and fails to precisely describe the relative size of chairs, as shown at the bottom row. However, after fine-tuning on our benchmark, LLaVA-1.5 accurately describes the curved shape of the banana and correctly identifies the two largest chairs. These improvements in visual question answering quality are observed across the testing set, demonstrating the effectiveness of our fine-tuning-based VH mitigation method.
Appendix C Prompts Used in VHTest
@@
Prompt for text description generation using a successful initial VH instance: [image] Question: [question] Multi-modal LLM (MLLM)’s response: [testing MLLM’s hallucinated response] Detected hallucination in the response: [detected hallucination in the response] Focus on the following elements: image, question, MLLM’s response, and the detected hallucination in the response. I’m trying to identify visual hallucinations in the MLLM associated with its visual process. Through the specific example, are there any general types of hallucination modes you notice the MLLM makes, or any visual features that MLLM fails to encode, ultimately leading to the errors in MLLM’s response? Try to give hallucination modes that are specific enough that someone could enable consistent reproduction of images and corresponding questions. Please try to include as many general hallucination modes as possible. These hallucination modes will be used later to generate images or videos. In your hallucination modes, please clearly explain why the this hallucination mode would cause difficulties for the vision encoder of MLLM to understand images related to this hallucination mode. I will further use this reason to precisely generate images that match the hallucination mode and are able to mislead the MLLM. Please encapsulate the essence of the examples provided, summarize as many as possible and stick to the examples.
@@
Prompt for text description generation using an unsuccessful initial VH instance: [image] Question: [question] A hypothetical response from the multi-modal LLM (MLLM): [a hypothetical hallucinated response] Detected hallucination in the hypothetical response: [detected hallucination in the hypothetical response] Focus on the following elements: image, question, MLLM’s hypothetical response, and detected hallucination in the hypothetical response. I’m trying to identify visual hallucinations in the MLLM associated with its visual process. Through the specific example, are there any general types of hallucination modes you notice the MLLM makes, or any visual features that MLLM fails to encode, ultimately leading to the errors in MLLM’s response? Try to give hallucination modes that are specific enough that someone could enable consistent reproduction of images and corresponding questions. Please try to include as many general hallucination modes as possible. These hallucination modes will be used later to generate images or videos. In your hallucination modes, please clearly explain why the this hallucination mode would cause difficulties for the vision encoder of MLLM to understand images related to this hallucination mode. I will further use this reason to precisely generate images that match the hallucination mode and are able to mislead the MLLM. Please encapsulate the essence of the examples provided, summarize as many as possible and stick to the examples.
@@
Prompt for text description integration: I want you to focus on the [VH mode]: [the definition of VH mode]. Try to summarize the [VH mode] text description in ONE paragraph less than 200 words to explain the causes of the hallucination based on the information below. [N text descriptions]
@@
Prompt for VH image generation: Generate an image that reflects the given hallucination mode. After your image generation, we will manually generate questions based on the image. Your MAIN GOAL is to ensure that the image aligns with the hallucination mode well so that querying the multi-modal LLM with the combination of manual questions and the image can effectively induce hallucinations. You will be evaluated on how well you actually perform. The generated image should ideally align with the hallucination mode, but there’s room for creativity. Be both creative and cautious. You can try to create image with different scenes and objects that align with the hallucination mode. Moreover, when you generate images, remember you are a very clever expert in exploiting the hallucination mode. For future debugging purposes, ensure that the generated images with our manual questions MUST cause multi-modal LLMs to provide incorrect responses. Hallucination Mode: [Text description of a VH mode]
@@
Prompt for an LLM to generate prompts for text-to-image generative models: Generate an image that reflects the given hallucination mode. After your image generation, we will manually generate questions based on the image. Your MAIN GOAL is to ensure that the image aligns with the hallucination mode well so that querying the multi-modal LLM with the combination of manual questions and the image can effectively induce hallucinations. You will be evaluated on how well you actually perform. The generated image should ideally align with the hallucination mode, but there’s room for creativity. Be both creative and cautious. You can try to create image with different scenes and objects that align with the hallucination mode. Moreover, when you generate images, remember you are a very clever expert in exploiting the hallucination mode. For future debugging purposes, ensure that the generated images with our manual questions MUST cause multi-modal LLMs to provide incorrect responses. Hallucination Mode: [Text description of a VH mode] As an LLM, DO NOT generate images, but generate prompts for text-to-image generative models aligning with the hallucination mode. I will use your generated prompts for text-to-image generative models later.
@@
Appendix D Question Templates
@@
Question templates for existence VH mode 1. Could you describe [object] in the picture? 2. Describe the [property] of [object] in the picture.
@@
Question templates for shape VH mode 1. Describe the shape(s) of [object] in the picture. 2. In the photo, is [object] depicted with a conventional, typical shape?
@@
Question templates for color VH mode 1. Describe the color(s) of [object] in the picture. 2. Describe [n] main color(s) of [object] in the picture. 3. Describe all the colors of [object] in the picture. Do not miss any color. 4. Describe the original color of [object] in the picture.
@@
Question templates for orientation VH mode 1. From the camera/viewer’s perspective, please choose the direction that is most proximate to the orientation in which [object] is facing in the picture. [Left, 45 degrees leftwards away from the camera/viewer, Away from the camera/viewer, 45 degrees rightwards away from the camera/viewer, Right, 45 degrees rightwards towards the camera/viewer, Towards the camera/viewer, 45 degrees leftwards towards the camera/viewer] 2. From the camera/viewer’s perspective, please choose the direction that is most proximate to the orientation in which [object] is facing in the picture. [Left, Left and upward within the pictorial plane, Upward within the pictorial plane, Right and upward within the pictorial plane, Right, Right and downward within the pictorial plane, Downward within the pictorial plane, Left and downward within the pictorial plane] 3. From the camera/viewer’s perspective, please choose the direction that is most proximate to the orientation in which [object] is facing in the picture. [Away from the camera/viewer, Towards the camera/viewer]
@@
Question templates for OCR VH mode 1. Read the (English) character(s)/word(s)/ caption [at WHERE] in the picture.
@@
Question templates for size VH mode 1. Could you identify and locate [object] with the largest/smallest size in the picture? 2. Could you identify and locate [object] with the tallest/shortest height in the picture? 3. Could you identify and locate [object] with the second largest/smallest size in the picture? 4. Could you identify and locate [object] with the largest/smallest and the second largest/smallest size in the picture? 5. Could you identify and locate [object] with the tallest/shortest and the second tallest/shortest height in the picture? 6. Could you identify and locate the 2/3/4 largest/smallest [object] in the picture? 7. Which one is larger/smaller/taller/shorter, [object A] or [object B]? 8. What is the largest [object] in the picture?
@@
Question templates for position VH mode 1. Is [object A] to the left/right of/on the top of/being placed on/above/under/inside/outside [object B]? 2. Which one is on the left/right/top, [object A] or [object B]? 3. Which one is closer to/further from the camera or viewer perspective, [object A] or [object B]? 4. Which one is positioned under/positioned above/closer to [reference object], [object A] or [object B]? 5. Regardless of the positional relationship of the actual scene in reality, from the camera’s perspective in the photo, is [object A] positioned above/below [object B]? 6. Is [object A] on and touching [object B]?
Question templates for counting VH mode 1. How many [object] are depicted/there/visible/can be seen [at WHERE] in the picture?
Appendix E Special Cases in Step III
@@
Prompt to find names of non-existent objects in a VH image in the existence VH mode [image] First, list all the names of objects in the picture. And then return some names of objects according to Requirement 1 or Requirement 2: Requirement 1: Objects you associate with the scene that should be there but are not actually there in the picture. Requirement 2: Objects that look similar to an object in the picture but do not actually exist in the picture.
@@
Appendix F Details of Fine-tuning Experiments
F.1 Hyperparamters
@@
F.2 Pre-processing of the Training Set
As mentioned in Liu et al. (2023), the short-form data overfit an MLLM behaviorally to short-form answers. We follow this work to do pre-processing on both counting and position VH modes. To be more specific, for counting VH mode instances, we use the prompt below in Figure 15 for an LLM (e.g., GPT-4) to modify every number reference answer into a reference sentence. As for position VH mode instances, we add the sentence “Answer the question using a single word or phrase.” after all the questions in every position VH instance.
Prompt for reference answer modification in counting VH mode Don’t talk nonsense. Turn the ground-truth numbers into extremely correct, extremely accurate, and only a little of diverse sentences based on the corresponding question. Sentences should be independent of each other, with each sentence occupying one line.